OpenStudy (anonymous):

Solve the proportion 2/x= √3+1/5

mathslover (mathslover):

@tamlin first of all welcome to openstudy . \[\huge{\frac{2}{x}=\sqrt{3}+\frac{1}{5}}\] is this the question ?

OpenStudy (anonymous):

except 3+1 is under a longer line together

Parth (parthkohli):

Aha \(\Rightarrow \Large {2 \over x} = {\sqrt{3 + 1} \over 5} \)

Parth (parthkohli):

\(\Rightarrow \Large{ 2 \over x} = {{\sqrt3 } + 1 \over 5} \) \(\Rightarrow 2 \times 5 = x(\sqrt3 + 1) \)

mathslover (mathslover):

ok : \[\huge{\frac{2}{x}=\sqrt{3}+\frac{1}{5}}\] \[\huge{\frac{2}{x}=\frac{(5\sqrt{3}+1)}{5}}\] \[\huge{2*5 = x* (5\sqrt{3}+1)}\]

Parth (parthkohli):

\(\Rightarrow 10 = x\sqrt3 + x\) Can you solve?

Parth (parthkohli):

You're welcome :)

mathslover (mathslover):

from my side too..

Parth (parthkohli):

\(\Rightarrow 10 - x = x\sqrt3\) \(\Rightarrow \Large {10 - x \over \sqrt3} = x\) \(\Rightarrow \Large{10\sqrt{3} - x\sqrt 3 \over 3} = x\) \(\Rightarrow 3x = 10\sqrt3 - x\sqrt3\) Can you solve now?

Parth (parthkohli):

\(\Rightarrow 3x = (10 - x)\sqrt3\) \(\Rightarrow \Large {3x \over \sqrt3} = 10 - x\) \(\Rightarrow \Large {3x\sqrt3 \over 3} = 10 - x\) \(\Rightarrow 3x\sqrt3 = 30 - 3x\) \(\Rightarrow 3x\sqrt 3 - 30 = -x\) \(\Rightarrow x = -3x\sqrt3 + 30\)